# Using shadow of black hole to examine the energy exchange between axion matter and rotating black hole

### v1

@@ Abstract
We find that a slowly rotating axion-modified black hole resulting from the backreaction of an axion field on a rotating Kerr black hole can have a D-shaped shadow as that for a highly counter-rotating Kerr black hole. This attributes to the fact that the energy exchange between the axion matter and the black hole influences the rotation of the black hole, so the black hole angular momentum first decreases to zero and then the BH starts to rotate to the opposite direction. Further increasing the coupling leads to a “human-face-like” shaped shadows and new lensings due to the chaotic scattering, which are novel and drastically different from Kerr black hole. Our analysis provides the first counterexample to that slowly rotating black hole has nearly circular shadow.
@@ I Introduction
In a realistic Universe the black holes (BHs) are rotating which are described by the Kerr metric Kerr:1963ud . For a rotating BH described by the Kerr metric, the description of the behaviour of matter is more complicated than in the non-rotating case. The metric in this case is only stationary, whilst the spherical symmetry has been replaced by axisymmetry. A further difficulty is the complexity of the coordinate systems for describing processes near a Kerr BH. The behaviour of matter in the neighbourhood of rotating BHs has been extensively studied. In the pioneering work, Bardeen, Press and Teukolsky had studied the physical processes around the BHs, together with their properties and interactions with the surrounding matter Bardeen:1972fi . To avoid the complexity in the perturbation analysis, in most cases it is assumed that the background metric is slowly rotating. Their method has essentially assumed that the non-rotating matter cancels out the “frame-dragging” effects of the BH rotation. In a sense, such a matter is co-rotating with the BH in such a way that the physical processes can be analyzed in their own frame. The authors of Bardeen:1972fi have used the method of locally non-rotating frames, which gives a clear picture of the process of energy extraction. The amount of such energy however, for realistic BHs, is not astrophysically significant.
In modified theories of gravity like the Chern-Simons (CS)-gravity theories with the action Jackiw:2003pm ; Zanelli:2005sa ; Alexander:2009tp ; Yunes:2009hc ; Yagi:2012ya
@@
where $b$ is a pseudoscalar denoting the axion matter field, $A$ is the coupling parameter of the axion field to the gravitational CS term $R_{CS}$ with dimension of length, and $\kappa^{2}=8\pi G$ with $G$ is the Newton’s gravitational constant in four dimensional spacetime. Note that we shall use natural units $G=c=1$. The CS modification is a total derivative, and thus innocuous unless pseudoscalar (axion-like) fields couple to it, in which case it deforms non-trivially the gravitational field of spinning black holes. Such deformations are described by a modified Kerr geometry. These Kerr-like BH solutions contain second-order curvature corrections (fourth-order in derivatives), and also characterise the low-energy-limit of string theory Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz . In the context of string theory, the axion fields are part of the excitation spectrum, and they are of two types Svrcek:2006yi . One, is the so-called string-model independent axion, which in (3+1)-dimensions is dual Duncan:1992vz ; Svrcek:2006yi to the field strength of the so-called Kalb-Ramond, spin-one, antisymmetric-tensor field, which plays the role of a totally antisymmetric torsion in the geometry. The second type of axions in string theory arise from compactification of the extra dimensions. Potential signatures of CS theories in the emission of gravitational waves, produced in the inspiral of stellar compact objects into massive black holes, both for intermediate and extreme mass ratios, have been investigated in Sopuerta:2009iy . Also in Canizares:2012is , the authors attempted to answer the question as to what extent the extreme mass ratio inspirals observations with a space based gravitational wave observatory like LISA might be able to distinguish between CS theories and General Relativity (GR).
@@
Rotating BHs have been studied extensively in such CS theories. First, they were found as solutions of second-order field equations for the metric stemming from string theories, in the presence of axions, in Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz , and more recently in Boulware:1985wk ; Gonzalez:2010vv . In particular, by extending non-trivially previous results Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz ; Yunes:2009hc ; Yagi:2012ya ; Myung:2020etf , slowly rotating BH solutions, in the presence of axion fields back reacting on the geometry, have been recently investigated Chatzifotis:2022mob . Its metric is
@@
$\begin{split}W(r)=\frac{2M}{r^{3}}-\frac{\gamma^{2}(189M^{2}+120Mr+70r^{2})}{14r^{8}}+O(A^{2n})~{},~{}~{}~{}~{}\gamma^{2}=\frac{A^{2}\kappa^{2}}{M^{4}}~{},\end{split}$ (3)
@@
where $M$, $\gamma$ and $a$ are the mass of BH, coupling parameter, and rotation parameter, respectively. The back reaction of the axionic field on the slowly-rotating background metric has been studied to arbitrary order in perturbation theory of an appropriate parameter, $\gamma$, proportional to the axion-CS interaction coupling. The expansion generated by the perturbation theory allows one to go arbitrarily close to the horizon of the hairy (with axion-hair Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz ) BH. One can thus study the behaviour of the axionic matter outside the horizon of the slowly-rotating Kerr-type metric and see how this matter distribution influences the Kerr-like BH itself. In the case of the Kerr BH, the total angular momentum is given by $Ma$ Carter:1968rr . In this respect we stress that in Chatzifotis:2022mob , the presence of the axion field coupled to the CS term, which back reacts on the Kerr BH background, implies that the total angular momentum consists of the contributions of both the axion field and the hairy Kerr-like BH itself, which arises as a consequence of the axionic back reaction on the geometry.
@@
In Chatzifotis:2022ene the study in Chatzifotis:2022mob of the axion-modified-Kerr BH system was extended by searching for potentially observable effects associated with bound trajectories of satellites in such backgrounds. As in Chatzifotis:2022mob a key role in the analysis will be played by the aforementioned perturbative parameter $\gamma$, which is proportional to the coupling of the axion field with the CS term and inversely proportional to the square of the total energy of the spacetime. This dimensionless parameter, rather than the coupling constant itself, that appears in the correction function of the modified Kerr-like BH solution and measures the strength of the back reaction on the geometry. The axion matter outside the horizon of the BH acquires an angular momentum in such a way that the total angular momentum of the system (BH and axionic matter) remains constant and equal to $Ma$. It was shown that as $\gamma$ increases, the BH angular momentum decreases and there exist a critical value $\gamma_{c}$ beyond which the BH angular momentum increases while the BH starts to rotate to the opposite direction. As such, the angular momentum of the BH horizon may in principle reach large values in magnitude, while the slowly rotating approximation remains valid with a total angular momentum equal to $Ma$. We understand this effect as a counterplay of two competing systems; the Kerr-like BH and the axionic matter rotating outside the horizon of the BH.
@@
In this work we study possible observational effects, like shadow and lensing Perlick:2021aok ; Cunha:2018acu ; Wang:2022kvg ; Lukes-Gerakopoulos:2010ipp , of the axion-modified-Kerr BH with the metric (2). The coupling between the axion and gravity modifies the equations of motions in such a way as to allow an exchange of energy between the axion field and gravity. We expect to probe the evidence of energy exchange by BH shadow, and provide alternative BH shadow profile, since nowadays is golden era to study various phenomenological deviations from GR BH but has potential to be detected by Event Horizon Telescope. Throughout all physical quantities are rescaled by $M$ so that we will work with $M=1$.
@@ II Shadow and Lensing by Axion induced Kerr-like black hole
$\mathcal{H}=-\frac{\partial S}{\partial\lambda}=\frac{1}{2}g_{\mu\nu}\frac{\partial S}{\partial x^{\mu}}\frac{\partial S}{\partial x^{\nu}}=0~{},$ (4)
@@
where $g_{\mu\nu}$ is the metric components of Kerr-like BH. Since the optical paths are reversible, the light rays reaching the observer can be considered to be emitted by the observer. These light rays have two endings, either returning to the light source or falling into the BH, where the light rays falling into the black hole correspond to the shadow of the BH.
@@
In Fig. 1, the images of Schwarzschild BH, Kerr BH $(a=0.999,-0.999)$, and Kerr-like BHs with different coupling parameters $\gamma$ are plotted. The observer is placed at $\theta_{o}=\pi/2$, the images have obvious symmetry, and the lensing effect can be clearly observed. The black area in the center is the shadow of the BH, and the white ring outside the shadow is the Einstein ring. The shadow of Schwarzschild BH is a perfect circle as can be seen in Fig. 1(b). Once we turn on the rotation in Kerr BH, $D$-shape will appear but it is faint, unless the spin $a$ is large enough, see for example $a=0.999$ (Fig. 1(c)) and $a=-0.999$ (Fig. 1(d)).
From the above figures it is easy to understand that for $a=0.2$ and $\gamma=0$, a tiny distortion exists in the left side. As we increase $\gamma$, the left distortion will become fainter because increasing $\gamma$ makes the BH angular momentum decrease but not yet change the rotating direction. There should be a critical $\gamma_{c}$ for which the angular momentum vanishes such that the shadow recover perfect circle. If we further increase $\gamma$ beyond $\gamma_{c}$, as shown in Chatzifotis:2022ene the BH angular momentum increases while it starts to rotate to the opposite direction, so the shadow distortion could appear in the other side because of the counter-rotating effect. This indeed explicitly happens in Fig. 1(h) which is similar to that in Kerr BH with $a=-0.999$ in Fig. 1(d). Therefore, starting with a slowly co-rotating BH with $a=0.2$ as we increase the coupling of the axion field to the CS term, because there is an exchange of energy between the axion field and the strong gravity, we observe a shadow which mimics that of a fast counter-rotating Kerr BH.
@@
Figure 3: Shadow boundary of the slowly rotating Kerr-like BHs extracted from Fig. 2.
In Fig. 2 and Fig. 3, the shadows and lensing effect of the slowly rotating Kerr-like BH with larger coupling parameters $\gamma$ are plotted. Comparing to the Kerr BH, we observe two novel structures of the shadow for the Kerr-like BH with $a=0.2$ : (i) The deformation of shadows of the slowly rotating Kerr-like BH with positive $a=0.2$ is on the same side of Kerr BH with negative $a$. Moreover, though the whole spacetime rotates slowly, the distortion of the shadow is significant. This means that non-trivial interaction between the axion and gravity could construct a slowly rotating spacetime as a whole, but has an internal structure in which a highly rotating event horizon is allowed, leading to highly rotating effects. (ii) Instead of standard $D$-shaped shadow for Kerr BH, here we observe “human-face-like” shaped shadow. This may provide a new attempt for us to use very large baseline interferometry (VLBI) observations EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022apq of the shadows to verify the existence of axion matter. Moreover, two disconnected crescent-shaped shadows appeared directly above and below the “human-face-like” shaped shadow, which is more explicit in Fig.3. Besides, the lensing effects, in particular in the vicinity of the distorted shadow, are novel and different from Kerr BH. As we comparing the lensing from Fig. 2(a)-Fig. 2(d), we detect that a regime with pink and blue lights will merge on the symmetric axial near the distorted shadow curve, as the coupling $\gamma$ increases. It means that in this Kerr-like BH, some photons can have more then one radial turning point, generically corresponding to chaotic motion Zelenka:2017aqn ; Destounis:2021mqv ; Destounis:2020kss , which distinguish from only one radial turning point in Kerr BH Wilkins:1972rs . The disconnected shadows and lensings show that the interaction between the axionic matter and gravity, and the distribution of axion significantly affect the chaotic scattering, so does the chaotic region.
@@ III Conclusion and discussion
Starting from an axion modified Kerr-like BH, we observe that a slowly rotating Kerr-like BH can produce a similar $D$-shaped shadow as that for a rapidly counter-rotating Kerr BH. These phenomena seems to be a natural consequence of the energy exchange between the axion field and BH such that a slowly rotating spacetime can also form a rapidly rotating event horizon. As the interplay becomes stronger, novel “human-face-like” shaped shadows and peculiar lensing phenomena, which significantly shift from those of Kerr BH, can be observed. Apparently, the rich lensing effects attributes to the chaotic scattering of the photons around the Kerr-like BH. An exhaustively studies on the bound orbit, light ring, and trapped trajectories Cunha:2016bjh could help us to obtain a deeper understanding of all those phenomenon, which shall be present elsewhere. In addition, it will be worthwhile to consider fast rotating Kerr-like BH to further study the effects among the fast rotation parameter $a$ and coupling parameter $\gamma$ on the shadows in the future.

### v3

# Using the shadow of a black hole to examine the energy exchange between axion matter and a rotating black hole

@@ Abstract
We find that a slowly rotating axion-modified black hole resulting from the backreaction of an axion field on a rotating Kerr black hole can have a D-shaped shadow as that for a highly counter-rotating Kerr black hole. This attributes to the fact that the energy exchange between the axion matter and the black hole influences the rotation of the black hole, so the black hole angular momentum first decreases to zero and then starts to rotate to the opposite direction. Further increasing the coupling leads to “human-face-like” shaped shadows and new lensing due to the chaotic scattering, which are novel and drastically different from those for Kerr black hole. Our analysis provides the first counterexample to that slowly rotating black hole has nearly circular shadow.
@@ I Introduction
In a realistic Universe, the black holes (BHs) are rotating which are described by the Kerr metric Kerr:1963ud . For a rotating BH described by the Kerr metric, it is more complicated to describe the behaviour of matter than in the non-rotating case. The metric in this case is only stationary, and the spherical symmetry has been replaced by axial symmetry. A further difficulty is the complexity of the coordinate systems for figuring out various processes near a Kerr BH. Till now, the behaviour of matter in the vicinity of rotating BHs has been extensively studied. In the pioneering work Bardeen:1972fi , the authors have studied the physical processes around the BHs, as well as their properties and interactions with the surrounding matters. In this scenario, one usually assumes that the background metric is slowly rotating to avoid the complexity in the perturbation analysis, the essence of which is to assume that the non-rotating matter cancels out the “frame-dragging” effects of the BH’s rotation. In a sense, such a matter is known to co-rotate with the BH such that the physical processes can be analyzed in its own frame. The authors of Bardeen:1972fi used the method of locally non-rotating frame, which could provide a clear picture to illustrate the process of energy extraction. However, the amount of such energy for realistic BHs is insignificant for astrophysics.
The modified theory of gravity like the Chern-Simons (CS)-gravity theory takes the action Jackiw:2003pm ; Zanelli:2005sa ; Alexander:2009tp ; Yunes:2009hc
@@
where $b$ is a pseudoscalar denoting the axion matter field, $A$ is the coupling parameter of the axion field to the gravitational CS term $R_{CS}$ with length dimension, and $\kappa^{2}=8\pi G$ with $G$ the Newton’s gravitational constant in four dimensional spacetime. Here, we shall use natural units $G=c=1$. It is noticed that the CS modification is a total derivative, and thus innocuous unless the axion-like pseudoscalar fields couple to it. The coupling gives a scalar field equation sourced by the Pontryagin density, while modifies the metric field equations into third-order derivatives, which may cause the full CS gravity not be a well-posted initial value problem Delsate:2014hba such that stable BH solutions are difficult to be found. But CS gravity is usually treated as an effective field theory characterising the low-energy-limit behaviors of string theory, thus, it is always treated to perturbatively deform from General Relativity (GR), reducing the order of the field equations Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz . In this case, the interaction could deform non-trivially the gravitational field of the spinning BHs, leading to a modified Kerr geometry (Kerr-like BH solution), which contains second-order curvature corrections. In string theory framework, the axion fields are part of the excitation spectrum, which in general have two types Svrcek:2006yi . One type is the so-called string-model independent axion. In (3+1)-dimensional case, this axion is dual Duncan:1992vz ; Svrcek:2006yi to the field strength of the Kalb-Ramond, spin-one, antisymmetric-tensor field, which plays the role of a totally antisymmetric torsion in the geometry. The other type of axion arises from the compactification of the extra dimensions in string theory. Potential signatures of CS theories in the observational gravitational waves (GW), emitted from the inspiral process of stellar compact objects into the massive BHs, with both intermediate and extreme mass ratios, have been explored in Sopuerta:2009iy . The authors of Canizares:2012is made a first attempt to answer the question that to some extent the observations from extreme mass ratio inspirals with a space-based gravitational wave observatory such as LISA might have potential to distinguish CS theories from GR. By simulating GW150914 with numerically binary BH signals in a CS theory, the leading order CS corrections to the merger waveform as well as the ringdown waveform and quasi-normal mode spectrum have been present in Okounkova:2019zjf . Furthermore, the possibility of detecting CS amplitude birefringence with GW detectors have been discussed in Yunes:2010yf ; Yagi:2017zhb ; Nojiri:2019nar ; Zhao:2019szi , which later been verified with real GW data from binary BH mergers Okounkova:2021xjv .
@@
In such CS theories, rotating BHs have been extensively studied. First, in Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz and recently in Boulware:1985wk ; Gonzalez:2010vv ; Yagi:2012ya , they were found as solutions to second-order field equations for the metric stemming from string theories with axions. Afterwards, focusing on some slowly rotating hairy BHs in this gravity, the BH perturbations Wagle:2023fwl , stable quasinormal spectrums Wagle:2021tam ; Srivastava:2021imr and superradiant instability Alexander:2022avt have been extensively explored. In particular, in Chatzifotis:2022mob , the authors extended the non-trivially previous results Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz ; Yunes:2009hc ; Myung:2020etf and investigated the slowly rotating BH solutions, in the presence of axion fields back reacting on the geometry, of which the metric is
@@
$\begin{split}W(r)=\frac{2M}{r^{3}}-\frac{\gamma^{2}(189M^{2}+120Mr+70r^{2})}{14r^{8}}+\mathcal{O}(A^{2n})~{},~{}~{}~{}~{}\gamma^{2}=\frac{A^{2}\kappa^{2}}{M^{4}}~{},\end{split}$ (3)
@@
where $M$, $\gamma$ and $a$ are the mass of BH, coupling parameter, and rotation parameter, respectively. The back reaction of the axion field on the slowly rotating metric has been studied to arbitrary order in the perturbation theory with an appropriate parameter, $\gamma$, which is proportional to the axion-CS coupling strength. It is explicit that instead of the coupling constant itself, the dimensionless parameter $\gamma$ appears in the correction function of the modified Kerr-like BH solution and measures the strength of the back reaction on the geometry. The expansion generated by this perturbation theory allows one to arbitrarily approach the horizon of the hairy BH. Thus, one could investigate the axionic matter’s behaviour around the horizon of the slowly rotating Kerr-like metric and check how the matter distribution would affect the geometry itself. Especially, it also means that the axion-hair is not short and can extend into the innermost light ring, which is crucial to the observations from BH shadow and gravitational lensing aspects Ghosh:2023kge . It is known that the total angular momentum for Kerr BH is given by $Ma$ Carter:1968rr . However, here the presence of the axion field coupled to the CS term, which back reacts on the original Kerr background, implies that the total angular momentum of the sector should consist of the contributions from both the axion field and the hairy Kerr-like BH itself, which arises because of the axionic matters back reaction to the geometry Chatzifotis:2022mob ; Chatzifotis:2022ene .
@@
Later, the study on this axion-modified-Kerr BH system was extended by seeking for potentially observable effects from the bound trajectories of satellites in such backgrounds Chatzifotis:2022ene , in which the aforementioned perturbative parameter $\gamma$ plays a key role. They found that the axion matters surrounding the BH horizon acquires angular momenta in such a way that the total angular momentum of the sector consisting of BH and axionic matter remains constant $Ma$. Consequently, the BH angular momentum decreases as $\gamma$ increases, and at a critical value $\gamma_{c}$, the BH starts to rotate to the opposite direction and the angular momentum increases when further increasing $\gamma$. Following the line, the angular momentum of the BH horizon can reach large values in magnitude, while the slowly rotating approximation is still valid with a total angular momentum that is equal to $Ma$, which was interpreted as the effects of a counterplay between two competing systems, i.e., the Kerr-like BH and the axionic matters rotating outside the BH horizon.
@@
In this work, we study possible observational effects, like shadow and lensing Perlick:2021aok ; Cunha:2018acu ; Wang:2022kvg ; Lukes-Gerakopoulos:2010ipp , of the axion-modified-Kerr BH with the metric (2). The coupling between the axion and gravity modifies the equations of motions in such a way as to allow an exchange of energy between the axion field and gravity. We expect to probe the evidence of energy exchange by BH shadow, and provide alternative BH shadow profile, since nowadays is golden era to study various phenomenological deviations from BH in GR but has potential to be detected by Event Horizon Telescope. Throughout the paper, all physical quantities are rescaled by $M$ so that we will work with $M=1$.
@@ II Shadow and Lensing by Axion induced Kerr-like black hole
$\mathcal{H}=-\frac{\partial S}{\partial\lambda}=\frac{1}{2}g^{\mu\nu}\frac{\partial S}{\partial x^{\mu}}\frac{\partial S}{\partial x^{\nu}}=0~{},$ (4)
@@
where $g_{\mu\nu}$ is the metric of Kerr-like BH. Since the optical paths are reversible, the light rays reaching the observer can be considered to be emitted by the observer. These light rays have two endings, either returning to the light source or falling into the BH, where the light rays falling into the BH correspond to the shadow of the BH.
@@
Figure 1: Shadow and Lensing by Schwarzschild BH, Kerr BHs and slowly rotating Kerr-like BHs. Here the inclination angle of the observer is fixed at $\theta_{o}=\pi/2$.
In Fig. 1, the images of Schwarzschild BH, Kerr BHs $(a=0.999,-0.999)$, and Kerr-like BHs with different coupling parameters $\gamma$ are plotted. The observer is placed at $\theta_{o}=\pi/2$. The black area in the center is the shadow of the BH which has obvious symmetry, and the white ring outside the shadow is the Einstein ring that can be clearly observed. The shadow of Schwarzschild BH is a perfect circle as can be seen in Fig. 1(b). Once we turn on the rotation in Kerr BH, $D$-shape will appear but it is faint, unless the spin $a$ is large enough, see for examples $a=0.999$ (Fig. 1(c)) and $a=-0.999$ (Fig. 1(d)).
From the figure it is easy to understand that for $a=0.2$ and $\gamma=0$, a tiny distortion exists in the left side of the central shadow because of the co-rotating. As we increase $\gamma$, the left distortion will become fainter because increasing $\gamma$ makes the BH angular momentum decrease but not yet change the rotating direction. There should be a critical $\gamma_{c}$ for which the angular momentum vanishes such that the shadow recovers perfect circle. If we further increase $\gamma$ beyond $\gamma_{c}$, as shown in Chatzifotis:2022ene the BH angular momentum increases while it starts to rotate to the opposite direction, so the shadow distortion could appear in the other side because of the counter-rotating effect. This indeed explicitly happens in Fig. 1(h) which is similar to that in Kerr BH with $a=-0.999$ in Fig. 1(d). Therefore, starting with a slowly co-rotating BH with $a=0.2$, as we increase the coupling of the axion field to the CS term, because there is an exchange of energy between the axion field and the strong gravity, we observe a shadow which mimics that of a fast counter-rotating Kerr BH.
@@
Figure 3: Shadow boundaries of the slowly rotating Kerr-like BHs extracted from Fig. 2.
In Fig. 2 and Fig. 3, the shadows and lensing effects of the slowly rotating Kerr-like BH with larger coupling parameters $\gamma$ are plotted. Comparing to the Kerr BH, we observe two novel structures of the shadow for the Kerr-like BH with $a=0.2$ : (i) The deformation of shadows of the slowly rotating Kerr-like BH with positive $a=0.2$ is on the same side of Kerr BH with negative $a$. Moreover, though the whole spacetime rotates slowly, the distortion of the shadow is significant. This means that non-trivial interaction between the axion and gravity could construct a slowly rotating spacetime as a whole. But it has an internal structure in which a highly rotating event horizon is allowed, leading to highly rotating effects. (ii) Instead of standard $D$-shaped shadow for Kerr BH, here we observe “human-face-like” shaped shadow. This may provide a new attempt for us to use very large baseline interferometry (VLBI) observations EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022apq of the shadows to verify the existence of axion matter. Moreover, two disconnected crescent-shaped shadows appeared directly above and below the “human-face-like” shaped shadow, which is more explicit in Fig. 3. It is noted that the critical shadow curves of other slowly rotating CS BHs have been studied in Amarilla:2010zq ; Meng:2023wgi , in which the novel shadow shapes in the current model were not observed. Besides, the lensing effects, in particular in the vicinity of the distorted shadow, are novel and different from Kerr BH. As we compare the lensing from Fig. 2(a) to Fig. 2(d), we detect that a regime with pink and blue lights will merge on the symmetric axial near the distorted shadow curve, as the coupling $\gamma$ increases. It means that in this Kerr-like BH, some photons can have more than one radial turning point, generically corresponding to chaotic motion Zelenka:2017aqn ; Destounis:2021mqv ; Destounis:2020kss , which distinguish from only one radial turning point in Kerr BH Wilkins:1972rs . The disconnected shadows and lensing show that the interaction between the axionic matter and gravity, and the distribution of axion significantly affect the chaotic scattering, so does the chaotic region.
@@ III Conclusion and discussion
Starting from an axion modified Kerr-like BH, we observe that a slowly rotating Kerr-like BH can produce a similar $D$-shaped shadow as that for a rapidly counter-rotating Kerr BH. These phenomena seem to be a natural consequence of the energy exchange between the axion field and BH such that a slowly rotating spacetime can also form a rapidly rotating event horizon. As the interplay becomes stronger, novel “human-face-like” shaped shadows and peculiar lensing phenomena, which significantly shift from those of Kerr BH, can be observed. Apparently, the rich lensing effects attribute to the chaotic scattering of the photons around the Kerr-like BH. An exhaustive studies on the bound orbits, light rings, and trapped trajectories Cunha:2016bjh could help us to obtain a deeper understanding of all those phenomena, which shall be present elsewhere. In addition, it will be worthwhile to consider fast rotating Kerr-like BH to further study the effects among the fast rotation parameter $a$ and coupling parameter $\gamma$ on the shadows in the future.